OpenStudy (anonymous):

factor: (x^2+9)^1/2 +6(x^2+9)^-1/2

OpenStudy (anonymous):

This might help :) http://www.purplemath.com/modules/factquad4.htm

OpenStudy (anonymous):

the farthest I get is (x^2+9)^-1/2 (6+(x^2+9))

OpenStudy (anonymous):

\[\left( x ^{2}+9 \right)^{1/2} +6\left( x ^{2}+9 \right)^{-1/2}\] First I believe this is the correct format for the expression

OpenStudy (anonymous):

yes it is @VeritasVosLiberabit

OpenStudy (anonymous):

What should be done first is to factor the term \[x ^{2}+9\]

OpenStudy (anonymous):

okay that's just (x-3)(x+3) right?

OpenStudy (anonymous):

yes that's right

OpenStudy (anonymous):

actually no it isn't

OpenStudy (anonymous):

You're right

OpenStudy (anonymous):

it's (x^2+9) which is not factorable

OpenStudy (anonymous):

i know that i need to find a common factor, which is (x^2+9)^-1/2

OpenStudy (anonymous):

yes, I have to think about this one, not sure how to solve it yet.

OpenStudy (anonymous):

\[\frac{ (x ^{2}+9)^{1/2}(x ^{2}+9)^{1/2} }{ (x ^{2}+9)^{1/2} }+\frac{ 6 }{ (x ^{2}+9)^{1/2} }\]

myininaya (myininaya):

\[u^\frac{1}{2}+6u^\frac{-1}{2} \\ u^\frac{-1}{2}(u^\frac{2}{2}+6) \\ u^\frac{-1}{2}(u+6)\]

myininaya (myininaya):

you can simplify that more

myininaya (myininaya):

you know by rewriting without the negative exponent

myininaya (myininaya):

@VeritasVosLiberabit 's way would have worked if he added 1/2 and 1/2 correctly :p

myininaya (myininaya):

remember if you have u^(1/2)*u^(1/2)=u^(1/2+1/2)=u^(2/2)=u

OpenStudy (anonymous):

@myininaya *facepalm* hehe. adding exponents is hard

OpenStudy (anonymous):

lol my final answer was right i just wasnt allowed to have negative exponents so i made 1/U (u+6)